A thermodynamic system undergoes a process in which its internal resistance decreases by 500J. Over the same time interval, 220J of work is done on the system. Find the energy transferred from it by heat.
1
Expert's answer
2021-02-17T11:09:57-0500

By the definition of the First Law of Thermodynamics, we have:


"\Delta E_{int}=Q+W,""Q=\Delta E_{int}-W=-500\ J-220\ J=-720\ J."